2 sin A / ( 1+ sin A+cos A)  = k , then prove that ,
(1+sinA+cosA)/(1+sinA) = k
 

Grade:11

1 Answers

Lab Bhattacharjee
8 years ago
HINT:$$(1+\sin A+\cos A)^2=(1+\sin A)^2+\cos^2A=\cdots=2(1+\sin A)$$HINT:$$(1+\sin A+\cos A)^2=(1+\sin A)^2+\cos^2A=\cdots=2(1+\sin A)$$ $$\\ \text{ and } \\(1+\sin A+\cos A)(1+\sin A-\cos A)=(1+\sin A)^2-\cos^2A=2\sin A(1+\sin A)$$